A sample of nitrogen gas occupies a volume of 1.8 L at 870 mm Hg and 5° C. The volume increases by 1.7 L and the temperature decreases to 45 K. What is the final pressure exerted on the gas in mm Hg? Round to two decimal places.


1
Expert's answer
2021-03-11T08:45:14-0500

Solution:

"\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}"


"\dfrac{870(1.8L)}{278K}=\dfrac{P_2(3.5L)}{45K}"



P2 = 72.42 mmHg